# Trajectory Optimization Under Stochastic Dynamics Leveraging Maximum Mean Discrepancy

### v1

@@ V Validation and Benchmarking
The objective of this section is to compare trajectory optimization with $r_{MMD}^{emp}$ as the risk cost vis-a-vis other popular alternatives. While doing so, we also expose the inner workings of our MMD-based approach.
@@ V-A Implementation Details
Figure 3: (3 Static obstacles) Ground-truth collision-rate achieved by minimum $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ trajectories (see (17)). We ensured that the respective optimal trajectories achieves $r_{MMD}^{emp}=0$, $r_{CVaR}^{emp}=0$. Thus, this result showcases the finite sample effectiveness of our MMD-based risk cost at low-sample regimes. We use the following values of the positive constants in 15a, 15b, 16a, 16b: Low Gaussian noise: $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.1,c_{\theta,2}=0.001$. High Gaussian noise: $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.15,c_{\theta,2}=0.001$. Low Beta noise: $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.001,c_{\theta,2}=0.001$. High Beta noise $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.0015,c_{\theta,2}=0.001$
Figure 4: (1 dynamic obstacle) Ground-truth collision-rate achieved by minimizing $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ (see (17)). The optimizer achieved $r_{MMD}^{emp}=0$, $r_{CVaR}^{emp}=0$ in all the scenarios. Thus, this result showcases the finite sample effectiveness of our MMD-based risk cost at low-sample regimes. We use the following values of the positive constants in 15a, 15b, 16a, 16b: Low Gaussian noise: $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.1,c_{\theta,2}=0.001$. High Gaussian noise: $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.15,c_{\theta,2}=0.001$. Low Beta noise $c_{a,1}=0.1,c_{a,2}=0.001,c_{\theta,1}=0.005,c_{\theta,2}=0.001$. High Beta noise: $c_{a,1}=0.15,c_{a,2}=0.001,c_{\theta,1}=0.0075,c_{\theta,2}=0.001$
(a) $r_{MMD}^{emp}$
(b) $r_{CVaR}^{emp}$
(c) $r_{MMD}^{emp}$
(d) $r_{CVaR}^{emp}$
Figure 6: ( MPC simulations in Carla T5 (a)-(b) and T10 (c)-(d) town. The red line is the route/reference-path. We show trajectories (green line) resulting from minimizing $r_{MMD}^{emp}$ (left column) and $r_{CVaR}^{emp}$ (right column) for $N=2$. The small number of samples of constraint residuals provide a very unreliable estimate of $r_{CVAR}^{emp}$ that changes sharply between consecutive MPC steps. As a result, the ego-vehicle is often unable to find any safe trajectory. As in the case shown, $r_{CVaR}^{emp}$ driven vehicle gets stuck behind the obstacle and eventually collides. In sharp contrast, $r_{MMD}^{emp}$ is able to regularly provide a consistent risk estimate for guiding the ego-vehicle. For town T5, the control noise for this particular experiment followed the Beta distribution with $c_{a,1}=0.05,c_{a,2}=0.4,c_{\theta,1}=0.05,c_{\theta,2}=0.01$ and for town T10, the control noise followed the Gaussian distribution with $c_{a,1}=0.3,c_{a,2}=0.3,c_{\theta,1}=0.3,c_{\theta,2}=0.01$
@@ V-A1 Benchmarking Environment
We perform benchmarking in trajectory-optimization and Model Predictive Control (MPC) setting. For the former, we randomly sample the initial states of the ego-vehicle, static obstacles and trajectories of dynamic obstacles, and use Alg.1 to compute an optimal trajectory. The latter involves constant re-planing using Alg.1 and the current state of the ego-vehicle.
For the benchmarking in MPC setting, we use high-fidelity simulator CARLA[12] to perform all our experiments. Each experiment requires the ego vehicle to maneuver from a start position to a goal position following a given route (reference path). An experiment is counted as successful if the ego vehicle makes a complete run without colliding with any of the obstacles. We perform experiments in two towns, namely, town 10(T10) and town 05(T5) . We consider a two-lane driving scenario with 8 and 10 static obstacles, respectively, in town 10 and town 05. We consider uncertainty in the dynamics of the ego vehicle, injected via perturbations of the nominal acceleration and steering control inputs.
@@ V-A2 Baselines
CVaR based Approach: This baseline is obtained by replacing $r_{MMD}^{emp}$ with a CVaR based risk defined in the following manner, in Alg.1.
@@
where $CVaR^{emp}$ is the empirical CVaR estimate computed using the constraint residual samples $\overline{h}({{}^{ij}}\mathbf{x})$ along the state-trajectory rollouts following the approach of [1]. Since, the underlying trajectory optimizer is same for minimizing both our proposed risk surrogate $r_{MMD}^{emp}$ and the baseline $r_{CVaR}^{emp}$, our benchmarking is designed to minimize optimizer bias and study the finite-sample performance of each risk estimate.
@@ V-A3 Metrics
In all our experiments, we fix the number of state-trajectory samples $N$ along which we evaluate $h(\mathbf{x})$ to estimate the risk. We recall, that our approach involves first computing $N^{2}$ state-trajectory rollouts and then choosing $N$ out of those for evaluating $h(\mathbf{x})$. This in turn limits the number of collision-checks that we need to perform. For CVaR baselines, the down-sampling process is not relevant and we compute only $N$ state-trajectory rollouts and evaluate $h(\mathbf{x})$ over them.
@@ V-B Benchmarking in Trajectory Optimization Setting
We created two sets of benchmarks with static and dynamic obstacles. For the former, we created 200 random scenarios and use two different noise models to perturb nominal dynamics. For the later, we had a single dynamic obstacle moving along a pre-defined trajectory and here again, we had two different noise models. The qualtiative trajectories in some of these scenarios are presented in the accompanying video. We use Alg.1 to compute minimum $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ trajectories in both static and dynamic obstacle benchmarks. Importantly, we ensured that the optimal trajectory obtained for each scenario in both the benchmarks had $r_{MMD}^{emp}=0$ and $r_{CVaR}^{emp}=0$. We then performed sampling around the optimal trajectory to compute the ground-truth collision-rate. To keep the analysis simple, this benchmarking did not enforce any risk cost associated with the lane violations.
The results are summarized in Fig.3,4 for different $N$. As can be seen that although both $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ are zero, the real collision-rates are very different. Our $r_{MMD}^{emp}$ results in lower collision-rate than $r_{CVaR}^{emp}$. For example, for low Gaussian noise and $N=4$, $r_{MMD}^{emp}$ trajectories provide almost two-times lower collision-rate than $r_{CVaR}^{emp}$ trajectories. Moreover, the collision-rate along $r_{MMD}^{emp}$ optimal trajectories shrink faster than they do along $r_{CVaR}^{emp}$ optimal trajectories. This result showcases the effectiveness of finite-sample estimate of our risk cost based on MMD vis-a-vis CVaR. This superior performance of $r_{MMD}^{emp}$ can be attributed to the fact the RKHS embedding is particularly effective in capturing the true distribution of arbitrary function of random variables with only a handful of samples (see Fig.1 in [6]). Our optimal reduced-set method (11a)-(LABEL:bi_level_inner_red_cost) further supercharges the capability of the RKHS embedding.
@@ V-C Benchmarking in MPC Setting using CARLA
We now evaluate the efficacy of $r_{MMD}^{emp}$ in a MPC setting where constant re-planning is done based on the current feedback of ego and the neighbouring vehicle state. In this benchmarking, we enforce a risk on the lane constraints as well. We also introduced Gaussian noise in the initial state.
Fig.6 presents a typical performance observed with $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ under Gaussian noise distribution. Due to extremely small $N$, the variance in $r_{CVaR}^{emp}$ estimation between consecutive MPC steps is large. As a result, the ego-vehicle often gets stuck behind the static obstacles and eventually collides. In sharp contrast, $r_{MMD}^{emp}$ estimates are consistently successful in guiding the ego-vehicle without collision. We demonstrate several such-scenarios in the supplementary video.
Table I presents the quantitative benchmarking in the MPC setting. We observed that re-planning can counter some of the effects of noise. Thus both $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ approaches show improvement over the performance we observed in a stand-alone planning setting of the previous sub-section. Nevertheless, $r_{MMD}^{emp}$ still outperformed $r_{CVaR}^{emp}$ in terms of collision-rate and achieved max and average speed. The lane violations of CVaR baseline was lower. However, this was because it had shorter runs due to collisions in several experiments. Table I also presents the results of the deterministic baseline. As can be seen, a noise ignorant approach is overconfident and moves too fast and consequently did not succeed in any of the simulation runs.
Table II repeats the benchmarking for higher noise level. Here, we only present results for our MMD-based approach and CVaR baseline. At higher noise, the performance (collision-rate, speed) of our $r_{MMD}^{emp}$ driven approach deteriorates a little, as compared to Table I. But, the degradation is far less than that observed for trajectories resulting from minimizing $r_{CVaR}^{emp}$.
TABLE I: $N=2$, Rollout horizon 40, 50 experiments
Gaussian $c_{a,1}=c_{\theta,1}=0.3$, Beta $c_{a,1}=c_{\theta,1}=0.01$
@@
Method/Town % Collisions (Gaussian/Beta) % Lane constr. (Gaussian/Beta) Avg. Speed(m/s) (Gaussian/Beta) Max. Speed(m/s) (Gaussian/Beta)
$r_{MMD}^{emp}$/T5 0/0 2.39/1.18 2.59/2.68 3.69/5.08
$r_{CVaR}^{emp}$/T5 7.69/0 2.25/1.08 2.12/2.49 3.04/4.61
DET/T5 100/100 0.03/0 5.25/5.27 8.72/8.15
$r_{MMD}^{emp}$/T10 0/0 0.47/0.99 3.42/3.63 5.94/5.5
$r_{CVaR}^{emp}$/T10 16.67/0 0.07/0.55 2.8/3.02 5.19/5.17
DET/T10 100/100 1.01/0.48 5.8/5.93 10.5/9.61
TABLE II: $N=2$, Rollout horizon 40, 50 experiments
Gaussian $c_{a,1}=c_{\theta,1}=0.3$, Beta $c_{a,1}=c_{\theta,1}=0.05$
@@
Method/Town % Collisions (Gaussian/Beta) % Lane constr. (Gaussian/Beta) Avg. Speed(m/s) (Gaussian/Beta) Max. Speed(m/s) (Gaussian/Beta)
$r_{MMD}^{emp}$/T5 0/0 2.7/3.5 2.59/2.02 4.06/4.07
$r_{CVaR}^{emp}$/T5 15/45 2.14/2.46 2.08/1.53 3.45/3.4
$r_{MMD}^{emp}$/T10 4/0 0.5/5.11 3.56/2 7.01/3.73
$r_{CVaR}^{emp}$/T10 17/4 0.1/3.23 2.73/1.56 4.55/4.07
@@ VI Conclusions and Future Work
Estimating state-dependent risk through forward simulation is a popular approach in motion planning. But computing rollouts can be expensive. More critically, evaluating constraints over the rollouts could be even more computationally prohibitive, for example, due to expensive collision checks. Thus, we presented a first such principled approach for reducing the number of constraint evaluations needed to estimate risk. Furthermore, our work can produce reliable results with a very few number of rollouts of system dynamics. Specifically, our work leveraged the solid mathematical foundations of RKHS embedding to propose a risk surrogate whose finite sample effectiveness can be improved by solving a bi-level optimization. We performed extensive simulations in both stand-alone trajectory optimization as well as MPC setting with a strong baseline based on CVaR. We showed that our MMD-based approach outperforms it in collision-rate and achieved average and maximum speed.

### v2

@@ V Validation and Benchmarking
This section compares trajectory optimization using $r_{MMD}^{emp}$ as the risk cost against other popular alternatives while also highlighting the inner workings of our MMD-based approach.
@@ V-A Implementation Details
TABLE I: (3 Static Obstacles) Low Gaussian noise: $c_{a,1}=0.1,c_{\theta,1}=0.1$. High Gaussian noise: $c_{a,1}=0.15,c_{\theta,1}=0.15$. Low Beta noise: $c_{a,1}=0.1,c_{\theta,1}=0.001$. High Beta noise: $c_{a,1}=0.15,c_{\theta,1}=0.0015$. $c_{a,2}=0.001$, $c_{\theta,2}=0.001$
[…]
N=2 N=4 N=6
Median Worst Median Worst Median Worst
Low Gaussian ${r}_{MMD}^{emp}$ 1.8 23.4 0.85 9.5 0.2 4
$r_{MMD-D}^{emp}$ 1.8 22.9 0.7 7.5 0.6 8.4
$r_{CVaR}^{emp}$ 4 32.5 1.25 16.4 1.35 10.4
High Gaussian $r_{MMD}^{emp}$ 3.9 31.4 2.8 15 1 9
$r_{MMD-D}^{emp}$ 4.2 32.2 1.8 10.9 0.9 9.3
$r_{CVaR}^{emp}$ 5 39.9 3.1 17.7 1.7 11.5
Low Beta $r_{MMD}^{emp}$ 0 5 0 1.7 0 1.5
$r_{MMD-D}^{emp}$ 0 4.5 0 1.6 0 1.5
$r_{CVaR}^{emp}$ 0.3 18.5 0 5.4 0 3
High Beta $r_{MMD}^{emp}$ 0.3 10 0 3.3 0 1.5
$r_{MMD-D}^{emp}$ 0.2 9.2 0 2.5 0 2.3
$r_{CVaR}^{emp}$ 0.5 21.8 0.4 8.1 0.25 6
TABLE II: (1 dynamic obstacle) Low Gaussian noise: $c_{a,1}=0.1,c_{\theta,1}=0.1$. High Gaussian noise: $c_{a,1}=0.15,c_{\theta,1}=0.15$. Low Beta noise: $c_{a,1}=0.1,c_{\theta,1}=0.005$. High Beta noise: $c_{a,1}=0.15,c_{\theta,1}=0.0075$. $c_{a,2}=0.001$, $c_{\theta,2}=0.001$
[…]
N=2 N=4 N=6
Median Worst Median Worst Median Worst
Low Gaussian $r_{MMD}^{emp}$ 4.1 41.3 0.25 9.2 0.55 9
$r_{MMD-D}^{emp}$ 4.4 35.1 1.35 13.2 0.8 8.7
$r_{CVaR}^{emp}$ 6.4 43.1 2.7 20.1 1.5 12.3
High Gaussian $r_{MMD}^{emp}$ 7.8 49.2 1.8 11.1 1.35 11.4
$r_{MMD-D}^{emp}$ 6.7 40.7 3 17.2 1.4 10.4
$r_{CVaR}^{emp}$ 9 46.3 3.5 19.7 1.6 11.3
Low Beta $r_{MMD}^{emp}$ 3.95 37.4 0.6 12 0.5 8.9
$r_{MMD-D}^{emp}$ 4.3 40.7 1.8 16.2 0.8 11.1
$r_{CVaR}^{emp}$ 8.5 55.6 3.5 28.3 1.6 12
High Beta $r_{MMD}^{emp}$ 5.1 35.2 0.65 12.9 1.4 15.2
$r_{MMD-D}^{emp}$ 13 78.9 5.85 39.5 2.4 18.4
$r_{CVaR}^{emp}$ 17.3 80.4 5.6 36.3 3.5 22.2
@@ V-A1 Benchmarking Environment
We benchmark in trajectory optimization and Model Predictive Control (MPC) settings. For trajectory optimization, we randomly sample initial states for the ego vehicle, static obstacles, and dynamic obstacle trajectories, using Alg.1 to compute the optimal trajectory. In the MPC setting, we continually re-plan with Alg.1 based on the ego vehicle’s current state.
Using the high-fidelity simulator CARLA[13], we conduct all MPC experiments where the ego vehicle maneuvers from a start to a goal position along a given reference path. An experiment is deemed successful if the vehicle completes the run without collisions. We test in town 10 (T10) and town 05 (T5), with a two-lane scenario featuring 8 obstacles in T10 and 10 in T5. We also introduce uncertainty in the ego vehicle’s dynamics through perturbations in nominal acceleration and steering control inputs.
@@ V-A2 Baselines
Ours with $r_{MMD-D}^{emp}$: This is similar to our main approach, except that the RKHS embedding and the MMD-based risk cost is constructed using the diagonal estimation derived in (14).
CVaR based Approach: This baseline replaces $r_{MMD}^{emp}$ with a CVaR based risk in Alg.1.
@@
where $CVaR^{emp}$ is the empirical CVaR estimate computed from constraint residual samples $\overline{h}({{}^{ij}}\mathbf{x})$ along the state-trajectory rollouts [1]. Since both $r_{MMD}^{emp}$ and $r_{CVaR}^{emp}$ use the same trajectory optimizer, our benchmarking minimizes optimizer bias to fairly evaluate their finite-sample performance.
@@ V-A3 Metrics
In all our experiments, we fix the number of state-trajectory samples $N$ along which we evaluate $h(\mathbf{x})$ to estimate the risk. We recall, that our main approach involves first computing $N^{2}$ state-trajectory rollouts and then choosing $N$ out of those for evaluating $h(\mathbf{x})$. This in turn limits the number of collision-checks that we need to perform. For CVaR baselines, the down-sampling process is not relevant and we compute only $N$ state-trajectory rollouts and evaluate $h(\mathbf{x})$ over them. Along similar lines, we also need only $N$ rollouts for the variant of our approach that uses diagonal estimation of RKHS embedding (14). It is worth re-iterating that our approach and all baselines have access to the same $N$ number of samples of constraint $h(\mathbf{x})$ evaluations.
@@ V-B Benchmarking in Trajectory Optimization Setting
We created two benchmark sets with static and dynamic obstacles. For static obstacles, we generated 200 random scenarios and applied two noise models to perturb nominal dynamics. For dynamic obstacles, a single obstacle followed a predefined trajectory, again with two noise models. Qualitative trajectories from these scenarios are shown in the accompanying video. Using Alg.1, we computed minimum $r_{MMD}^{emp}$, $r_{MMD-D}^{emp}$ (see (14)), and $r_{CVaR}^{emp}$ trajectories for both benchmarks. Crucially, each optimal trajectory satisfied $r_{MMD}^{emp}=0$, $r_{MMD-D}^{emp}=0$, and $r_{CVaR}^{emp}=0$. We then sampled around these optimal trajectories to compute the ground-truth collision rate. To simplify the analysis, this benchmarking did not impose any risk cost for lane violations.
MMD-vs-CVaR
The results in Tables I-II for different $N$ show that while $r_{MMD}^{emp}$, $r_{MMD-D}^{emp}$, and $r_{CVaR}^{emp}$ are all zero, actual collision rates vary significantly. Trajectories optimized using $r_{MMD}^{emp}$ or $r_{MMD-D}^{emp}$ consistently outperform those from $r_{CVaR}^{emp}$ across all $N$. For instance, in Table I (static obstacles, high-beta noise, $N=6$), $r_{MMD}^{emp}$ results in a worst-case collision rate that is twice as low as $r_{CVaR}^{emp}$, while $r_{MMD-D}^{emp}$ provides a smaller improvement. A similar trend appears in Table II for dynamic obstacles—under low-beta noise and $N=6$, MMD-based approaches achieve a median collision rate two to three times lower than the CVaR baseline.
The above result showcases the effectiveness of finite-sample estimate of our risk cost based on MMD vis-a-vis CVaR. This superior performance of $r_{MMD}^{emp}$ can be attributed to the fact the RKHS embedding is particularly effective in capturing the true distribution of arbitrary function of random variables with only a handful of samples (see Fig.1 in [7]). Our optimal reduced-set method (11a)-(LABEL:bi_level_inner_red_cost) further supercharges the capability of the RKHS embedding.
$r_{MMD}^{emp}$ vs $r_{MMD-D}^{emp}$
Table I-II shows that $r_{MMD-D}^{emp}$ which is based on diagonal estimation of RKHS embedding (recall (14)) can provide similar performance as our main approach based on $r_{MMD}^{emp}$. This is particularly true for low sample regime of $N=2$, as in this case, the reduced-set distillation will not add much value [^2] For $N=2$, we start with $N^{2}=4$ samples and down-sample to $N=2$; a difference of just two samples.. The performance difference is however more important at $N=6$, where $r_{MMD}^{emp}$ clearly shows improved performance.
@@ V-C Benchmarking in MPC Setting using CARLA
We now evaluate the efficacy of $r_{MMD}^{emp}$ in a MPC setting where constant re-planning is done based on the current feedback of ego and the neighboring vehicle state. In this benchmarking, we enforce a risk on the lane constraints as well. We also introduced Gaussian noise in the initial state. The qualitative results are presented in the accompanying video.
Table III presents the quantitative benchmarking in the MPC setting, where we also present results obtained with $r_{MMD-D}^{emp}$. We observed that re-planning can counter some of the effects of noise. All approaches improved over the stand-alone planning setting from the previous subsection. However, $r_{MMD}^{emp}$ and $r_{MMD-D}^{emp}$ outperformed $r_{CVaR}^{emp}$ in collision rate and achieved higher max and average speeds. Among them, $r_{MMD}^{emp}$ was superior in all benchmarks except one. The lane violations of CVaR baseline was lower. However, this was due to shorter runs caused by collisions in several experiments. Table III also includes the deterministic baseline, which, being noise-ignorant, was overconfident, moved too fast, and failed in all simulation runs.
Table IV repeats the benchmarking for higher noise level. Here, we only present results for our MMD-based approaches and CVaR baseline. At higher noise, the performance of $r_{MMD}^{emp}$ and $r_{MMD-D}^{emp}$ declines slightly (collision rate, speed) compared to Table III, but the degradation is significantly lower than in $r_{CVaR}^{emp}$-optimized trajectories.
TABLE III: $N=2$, Rollout horizon 40, 50 experiments.
Gaussian noise: $c_{a,1}=c_{\theta,1}=0.3$, Beta noise: $c_{a,1}=c_{\theta,1}=0.01$.
@@
Method Town % Collisions % Lane Constr. Viol. Avg. Speed (m/s) Max. Speed (m/s)
Gaussian Beta Gaussian Beta Gaussian Beta Gaussian Beta
$r_{MMD}^{emp}$ T5 0 0 2.39 1.18 2.59 2.68 3.69 5.08
$r_{MMD-D}^{emp}$ T5 0 2.63 0 1.23 2.05 2.72 3.26 4.79
$r_{CVaR}^{emp}$ T5 7.69 0 2.25 1.08 2.12 2.49 3.04 4.61
DET T5 100 100 0.03 0 5.25 5.27 8.72 8.15
$r_{MMD}^{emp}$ T10 0 0 0.47 0.99 3.42 3.63 5.94 5.5
$r_{MMD-D}^{emp}$ T10 0 0 0.58 1.3 3.61 3.91 6.99 6.44
$r_{CVaR}^{emp}$ T10 16.67 0 0.07 0.55 2.8 3.02 5.19 5.17
DET T10 100 100 1.01 0.48 5.8 5.93 10.5 9.61
TABLE IV: $N=2$, Rollout horizon 40, 50 experiments.
Gaussian noise: $c_{a,1}=c_{\theta,1}=0.3$, Beta noise: $c_{a,1}=c_{\theta,1}=0.05$.
@@
Method Town % Collisions % Lane Constr. Viol. Avg. Speed (m/s) Max. Speed (m/s)
Gaussian Beta Gaussian Beta Gaussian Beta Gaussian Beta
$r_{MMD}^{emp}$ T5 0 0 2.7 3.5 2.59 2.02 4.06 4.07
$r_{MMD-D}^{emp}$ T5 3.28 3.33 2.09 7.9 2.1 2.1 3.7 4.07
$r_{CVaR}^{emp}$ T5 15 45 2.14 2.46 2.08 1.53 3.45 3.4
$r_{MMD}^{emp}$ T10 4 0 0.5 5.11 3.56 2 7.01 3.73
$r_{MMD-D}^{emp}$ T10 2.44 0 0.94 8.4 3.55 2.29 7.17 4.24
$r_{CVaR}^{emp}$ T10 17 4 0.1 3.23 2.73 1.56 4.55 4.07
V-D Why MMD approaches performed better than CVaR?
CVaR captures tail risk but requires quantile selection and more samples for stability. In contrast, $r_{MMD}^{emp}$ uses RKHS embedding for richer statistical information, ensuring more accurate risk estimates especially in low-data regimes. Its optimal reduced-set selection further enhances sample efficiency, unlike CVaR.
V-E Computational Aspects
The $r^{emp}_{CVaR}$ baseline had the lowest computation time (0.1s per MPC step), while $r_{MMD}^{emp}$ took 0.18s due to the overhead from reduced-set computation. A simple way to reduce this is by running Alg.1 for fewer iterations, as shown in the accompanying video. Notably, the simplified $r_{MMD-D}^{emp}$ ran as fast as $r^{emp}_{CVaR}$ but performed significantly better, only slightly worse than $r_{MMD}^{emp}$ optimization. This trend highlights how our MMD-based methods balance performance and re-planning flexibility within an MPC setup. A detailed run-time comparison is provided in the video.
@@ VI Conclusions and Future Work
Estimating state-dependent risk through forward simulation is a popular approach in motion planning. But computing rollouts can be expensive. More critically, evaluating constraints over the rollouts could be even more computationally prohibitive, for example, due to expensive collision checks. Thus, we presented a first such principled approach for reducing the number of constraint evaluations needed to estimate risk. Furthermore, our work can produce reliable results with a very few number of rollouts of system dynamics. Specifically, we leveraged the solid mathematical foundations of RKHS embedding to propose a risk surrogate whose finite sample effectiveness can be further improved by computing the so-called reduced-set. We performed extensive simulations in both stand-alone trajectory optimization as well as MPC setting with a strong baseline based on CVaR. We showed that our MMD-based approach outperforms it in collision-rate and achieved average and maximum speed.